NOAA SATOPS Morning Report: Thursday, February 14, 2013


SATOPS Morning Report: Thursday, February 14, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 045

GOES
GOES-13 DOY044 The following images were lost or degraded due to a WCDA antenna problem: 21:45Z NHemi, 22:01z CONUS, 2146Z SOUNDER.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

SNPP
All SNPP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
ESPC DID NOT RECEIVE THE MET-10 IMAGES: 1445Z, 1500Z AND 1530Z. THIS WAS DUE TO UNSCHEDULED SYSTEM MAINTENANCE TO CORRECT MET-10 ISSUES. WALLOPS CDA ADVISED THEY WILL CONTINUE TO TROUBLESHOOT THESE ISSUES THE NEXT TWO DAYS.

DUE TO THE EAST GINI L1 MONITORING CRASHING WHICH CAUSED THE PROCESSING TO STOP, THE FOLLOWING GINI/AWIPS IMAGES WERE LOST ON THE EAST NPORT GINI. THE 1815Z, 1830Z ECONUS, 1745Z, 1815Z,1830Z PRNAT 1815Z ,1830Z PRREG 1745Z, 1815Z NHEMI AND THE1745Z 1815Z SUPERN. ESPC STOPPED AND RESTARTED THE REMAPPER AND DATA RESUMED PROCESSING.

DUE TO ANTENNA ISSUES AT WCDA FROM 2148Z UNTIL 2202Z, THE FOLLOWING PRODUCTS WERE LOST OR DEGRADED: GOES EAST - 2145Z NHEMI, 2146Z SOUNDER, 2202Z CONUS AND EAST LRIT DCSRB.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.